OpenStudy (anonymous):

This problem must die >.< but i have to do it, so help is loved :3

OpenStudy (anonymous):

\[9/\sqrt{20x}\]

OpenStudy (anonymous):

the / is a fraction. The problem says to 'rationalize the denominator'

OpenStudy (anonymous):

so rationalize... rationalize just means get rid of the square n the bottom

OpenStudy (anonymous):

so just multiply the top and bottom by sqrt(20x) to get just 20x on the bottom

OpenStudy (anonymous):

\[\frac{9}{\sqrt{20x}}*\frac{\sqrt{20x}}{\sqrt{20x}}=\frac{9\sqrt{20x}}{20x}\]

OpenStudy (anonymous):

\[=9\sqrt{4(5x)}=18\sqrt{5x}\]

OpenStudy (anonymous):

\[\frac{9\sqrt{5x}}{10x}\]

OpenStudy (anonymous):

Thanks :D